# FORGE: Forming Semantic Identifiers for Generative Retrieval in Industrial Datasets

### v1

@@ Abstract
Semantic identifiers (SIDs) have gained increasing attention in generative retrieval (GR) due to their meaningful semantic discriminability. However, current research on SIDs faces three main challenges: (1) the absence of large-scale public datasets with multimodal features, (2) limited investigation into optimization strategies for SID generation, which typically rely on costly GR training for evaluation, and (3) slow online convergence in industrial deployment. To address these challenges, we propose FORGE, a comprehensive benchmark for FOrming semantic identifieR in Generative rEtrieval with industrial datasets. Specifically, FORGE is equipped with a dataset comprising 14 billion user interactions and multimodal features of 250 million items sampled from Taobao, one of the biggest e-commerce platforms in China. Leveraging this dataset, FORGE explores several optimizations to enhance the SID construction and validates their effectiveness via offline experiments across different settings and tasks. Further online analysis conducted on our platform, which serves over 300 million users daily, reveals a 0.35% increase in transaction count, highlighting the practical impact of our method. Regarding the expensive SID validation accompanied by the full training of GRs, we propose two novel metrics of SID that correlate positively with recommendation performance, enabling convenient evaluations without any GR training. For real-world applications, FORGE introduces an offline pretraining schema that reduces online convergence by half. The code and data are available at https://github.com/selous123/al_sid.
@@ 1 Introduction
ii) The semantic identifiers shown in Figure 1(b) represent each item with multi-level codewords (e.g., <><><>
), where retrieval models would generate the top-K item identifiers as candidates with continuous beam search (li2025bbqrec, ; zheng2025universal, ). The construction of codewords within SID is based on multimodal data (e.g., text and images), enabling knowledge sharing among similar items. Furthermore, with hierarchical codewords, SIDs could compactly encode an exponentially large item space (e.g., a 3-level SID with 8192 tokens per level could represent items using 24576 tokens). Limited tokens also make it practical to include all tokens in the calculation for both training and serving (rajput2023recommender, ).
@@
We conduct extensive offline experiments to validate the effectiveness of our optimizations under different settings. Subsequent 0.35% online increase in transaction count within the "Guess You Like" section of Taobao further strengthens our credibility. Serving as an industrial work, FORGE discovers an offline pretraining approach to accelerate the convergence of new SIDs in production.
@@ 4.3 Analysis about the Generalization of SID Optimization (RQ3)
The left panel of Table 4 presents the performance of t5-base with an encoder-decoder architecture. While this architecture differs from the decoder-only setup we used in previous experiments, our proposed optimizations still yield substantial improvements across different architectural paradigms. The consistent performance gain over the base model for both types of SIDs highlights our broad applicability.
@@
The improvement through online experiments over 7 days in the Guess You Like Section of Taobao is shown in Table 6 to serve as strong evidence for our optimization, where our optimization improves the PVR and HR by 8.93% and 10.02% compared to base. Moreover, replacing base with ours ultimately results in a 0.35% absolute increase in user transactions.

### v2

@@ Abstract
Semantic identifiers (SIDs) have gained increasing attention in generative retrieval (GR) due to their meaningful semantic discriminability. However, current research on SIDs faces three main challenges: (1) the absence of large-scale public datasets with multimodal features, (2) limited investigation into optimization strategies for SID generation, which typically rely on costly GR training for evaluation, and (3) slow online convergence in industrial deployment. To address these challenges, we propose FORGE, a comprehensive benchmark for FOrming semantic identifieR in Generative rEtrieval with industrial datasets. Specifically, FORGE is equipped with a dataset comprising 14 billion user interactions and multimodal features of 250 million items sampled from Taobao, one of the biggest e-commerce platforms in China. Leveraging this dataset, FORGE explores several optimizations to enhance the SID construction and validates their effectiveness via offline experiments across different settings and tasks. Further online analysis conducted on the "Guess You Like" section of Taobao’s homepage shows a 0.35% increase in transaction count, highlighting the practical impact of our method. Regarding the expensive SID validation accompanied by the full training of GRs, we propose two novel metrics of SID that correlate positively with recommendation performance, enabling convenient evaluations without any GR training. For real-world applications, FORGE introduces an offline pretraining schema that reduces online convergence by half. The code and data are available at https://github.com/selous123/al_sid.
@@ 1 Introduction
ii) The semantic identifiers shown in Figure 1(b) represent items with multi-level codewords (e.g., <><><>
), and retrieval models would generate the top-K item identifiers as candidates with continuous beam search (li2025bbqrec, ; zheng2025universal, ). The construction of codewords within SID is based on multimodal data, enabling knowledge sharing among similar items. Furthermore, with hierarchical codewords, SIDs could compactly encode an exponentially large item space (e.g., a 3-level SID with 8192 tokens per level could represent items using 24576 tokens). The small number of tokens also makes it computationally feasible to include all of them in the calculations during training (rajput2023recommender, ).
@@
We conduct extensive offline experiments to validate the effectiveness of our optimizations under different settings. Subsequent 0.35% online increase in transaction count within the "Guess You Like" section of Taobao further strengthens our credibility. As for the industrial serving, FORGE discovers an offline pretraining approach to accelerate the convergence of new SIDs in production.
@@ 4.3 Analysis about the Generalization of SID Optimization (RQ3)
The left panel of Table 4 presents the performance of t5-base (raffel2020exploring, ) with an encoder-decoder architecture. While this architecture differs from the decoder-only setup we used in previous experiments, our proposed optimizations still yield substantial improvements across different architectural paradigms. The consistent performance gain over the base model for both types of SIDs highlights our broad applicability.
@@
The improvement from a 7-day online experiment conducted on the Guess You Like section of Taobao’s homepage is presented in Table 6, serving as strong evidence of our optimizations, where our optimization improves the PVR and HR by 8.93% and 10.02% compared to base. Moreover, replacing base with ours ultimately results in a 0.35% absolute increase in user transactions.